Star Videos, Inc., produces short musical videos for sale to retail outlets. The company’s balance sheet accounts as of January 1 are given below.

Star Videos, Inc.
Balance Sheet
January 1
Assets
Cash $ 90,200
Accounts receivable 96,200
Inventories:
Raw materials (film, costumes) $ 20,200
Videos in process 56,800
Finished videos awaiting sale 81,800 158,800
Prepaid insurance 12,700
Studio and equipment (net) 542,000
Total assets $ 899,900
Liabilities and Stockholders’ Equity
Accounts payable $ 156,000
Retained earnings 743,900
Total liabilities and stockholders’ equity $ 899,900

Because the videos differ in length and in complexity of production, the company uses a job-order costing system to determine the cost of each video produced. Studio (manufacturing) overhead is charged to videos on the basis of camera-hours of activity. The company’s predetermined overhead rate for the year ($40 per camera-hour) is based on a cost formula that estimated $280,000 in manufacturing overhead for an estimated allocation base of 7,000 camera-hours. Any underapplied or overapplied overhead is closed to cost of goods sold. The following transactions were recorded for the year:

  1. Film, costumes, and similar raw materials purchased on account, $193,500.
  2. Film, costumes, and other raw materials issued to production, $207,500 (85% of this material was considered direct to the videos in production, and the other 15% was considered indirect).
  3. Utility costs incurred (on account) in the production studio, $91,800.
  4. Depreciation recorded on the studio, cameras, and other equipment, $91,600. Three-fourths of this depreciation related to actual production of the videos, and the remainder related to equipment used in marketing and administration.
  5. Advertising expense incurred (on account), $173,500.
  6. Salaries and wages paid in cash as follows:
Direct labor (actors and directors) $ 97,600
Indirect labor (carpenters to build sets, costume designers, and so forth) $ 88,500
Administrative salaries $ 95,400
  1. Prepaid insurance expired during the year, $11,500 (70% related to production of videos, and 30% related to marketing and administrative activities).
  2. Miscellaneous marketing and administrative expenses incurred (on account), $14,450.
  3. Studio (manufacturing) overhead was applied to videos in production. The company recorded 7,250 camera-hours of activity during the year.
  4. Videos that cost $520,000 to produce according to their job cost sheets were transferred to the finished videos warehouse to await sale and shipment.
  5. Sales for the year totaled $956,000 and were all on account.
  6. The total cost to produce the videos that were sold according to their job cost sheets was $561,850.
  7. Collections from customers during the year totaled $906,000.
  8. Payments to suppliers on account during the year, $563,000.
  9. Underapplied or overapplied overhead $__?__.

Required:

1. Prepare a transaction analysis that records all of the above transactions. Calculate the ending balances at December 31 for all balance sheet accounts.

2. Prepare a schedule of cost of goods manufactured for the year.

3. Prepare a schedule of cost of goods sold for the year.

4. Prepare an income statement for the year.
Answer #1

1.

Star Videos, Inc.
Transaction Analysis
For the Year Ended December 31
Cash Accounts Receivable Raw Materials Videos in Process Finished Videos Manufacturing Overhead Prepaid Insurance Studio & Equipment (net) = Accounts Payable Retained Earnings
Beginning balance @ 1/1 90200 96200 20200 56800 81800 0 12700 542000 = 156000 743900
a) Raw material purchases 193500 = 193500
b) Raw materials used -207500 176375 31125 =
c) Utility costs 91800 = 91800
d) Depreciation charges 68700 -91600 = -22900
e) Advertising = 173500 -173500
f) Salaries & wages -281500 97600 88500 = -95400
g) Prepaid insurance 8050 -11500 = -3450
h) Miscellaneous marketing = 14450 -14450
i) Applied overhead 290000 -290000 =
j) Transfer completed videos to finished goods -520000 520000 =
k) Sales 956000 = 956000
l) Transfer finished goods to cost of goods sold -561850 = -561850
m) Cash collections from customers 906000 -906000 =
n) Payment to suppliers -563000 = -563000
o) Overapplied overhead 1825 = 1825
Ending balances @12/31 151700 146200 6200 100775 39950 0 1200 450400 = 66250 830175

2.

Star Videos, Inc.
Schedule of Cost of Goods Manufactured
For the Year Ended December 31
Beginning raw materials inventory $ 20200
Add: Raw material purchases 193500
Raw materials available for use 213700
Less: Ending raw materials inventory 6200
Raw materials used 207500
Less: Indirect materials 31125 176375
Direct labor 97600
Manufacturing overhead 290000
Total manufacturing costs 563975
Add: Beginning videos in process inventory 56800
Total videos in process inventory 620775
Less: Ending videos in process inventory 100775
Cost of goods manufactured 520000

3.

Star Videos, Inc.
Schedule of Cost of Goods Sold
For the Year Ended December 31
Cost of goods manufactured 520000
Add: Beginning finished goods inventory 81800
Cost of goods available for sale 601800
Less: Ending finished goods inventory 39950
Cost of goods sold 561850
Less: Overapplied overhead 1825
Adjusted cost of goods sold $ 560025

4.

Star Videos, Inc.
Income Statement
For the Year Ended December 31
Sales revenue 956000
Cost of goods sold 560025
Gross profit 395975
Operating expenses 309700
Net income $ 86275